# Improving Adversarial Robust Fairness via Anti-Bias Soft Label Distillation

### v1

@@ 4.1 Overall Framework
Based on the above finding, we can find that adjusting the class-wise smoothness degree of soft labels can be regarded as a potential way to obtain robust fairness. Then we consider introducing this ideology into Knowledge Distillation (KD), which has been proven to be an effective method to improve the robustness of small models [42, 43, 40, 12, 41]. Since the core idea of KD is to use the teacher’s soft labels to guide the student’s optimization process, we can adjust the class-wise smoothness degree of soft labels and obtain the student with both strong robustness and fairness.
@@ 4.2 Re-temperate Teacher’s Soft Labels
In order to adjust the class-wise smoothness degree of soft labels in ARD, two options exist: one is to use student feedback to update the teacher parameter in the process of optimizing students, but this option requires retraining the teacher model, which may bring the pretty optimization difficulty and computational overhead; the other is to directly adjust the smoothness degree of soft labels for different classes. Inspired by [41], we apply the temperature as a means of directly controlling the smoothness degree of soft labels during the training process. Here, we provide Theorem 4.1 about the relationship between the teacher’s temperature and the student’s class-wise error risk gap.
Theorem 4.1
@@
The proof of Theorem 4.1 can be found in Appendix B. From the actual optimization perspective, the Theorem 4.1 demonstrates that the different temperatures can influence the student robust fairness: when student’s error risk of $k$-th class is larger than the average error risk, we think that this type of class is relatively hard compared with others, then the teacher temperature for $k$-th class will reduce and the smoothness degree of soft labels will be sharper, and the optimization gap between teacher distribution and student distribution in $k$-th class will corresponding increase, leading to stronger learning intensity for $k$-th class and final reduce the student’s class-wise optimization error risk gap.
@@
Meanwhile, according to [37], both clean examples and adversarial examples exist fairness problem and can affect each other, so it is necessary to achieve fairness for both types of data. Since clean examples and adversarial examples of the same classes may have different error risks during the training process, it is unreasonable to use the same set of class temperatures to adjust both clean examples and adversarial examples. Here we simultaneously adjust the student’s clean error risk $R(f_{s}(x_{k}))$ and the student’s robust error risk $R(f_{s}(\tilde{x}_{k}))$, in other words, we apply two different sets of teacher temperatures: $\tau_{k}^{t}$ and $\tilde{\tau}_{k}^{t}$, for the adjustment of the teacher’s soft labels for clean examples and adversarial examples, respectively.
@@
1:Student $f_{s}$ with random initial weight $\theta_{s}$ and temperature $\tau_{s}$, pretrained robust teacher $f_{t}$, the initial temperature $\tau_{y}^{t}$ and $\tilde{\tau}_{y}^{t}$ for the teacher’s soft labels for clean examples and adversarial examples, where $y=\{1,\dots,C\}$, the train dataset $\mathcal{D}$, the maximum perturbation scale $\epsilon$, the max training epochs $max$-$epoch$.
@@
4:          $R(f_{s}(\tilde{x}_{k}))=0,~{}R(f_{s}(x_{k}))=0$.
@@
7:          $\tilde{x}=\mathop{argmax}\limits_{||\tilde{x}-x||\leq\epsilon}KL(f_{s}(\tilde{x};\tau^{s}),f_{t}^{{}^{\prime}}(x;\tilde{\tau}_{y}^{t}))$.
8:          $\theta_{s}=\theta_{s}-\eta\cdot\nabla_{\theta}\mathcal{L}_{absld}(\tilde{x},x;f_{s},f_{t}^{{}^{\prime}})$.
9:          $R(f_{s}(\tilde{x}_{y}))=R(f_{s}(\tilde{x}_{y}))+CE(f_{s}(\tilde{x}),{y})$.
10:          $R(f_{s}(x_{y}))=R(f_{s}(x_{y}))+CE(f_{s}({x}),{y})$.
@@
13:          Update $\tilde{\tau}_{k}^{t}$ and ${\tau}_{k}^{t}$ according to Eq.(11).
@@
$\mathcal{L}_{absld}$ $(\tilde{x},x;f_{s},f_{t}^{{}^{\prime}})=\alpha\frac{1}{C}\sum_{i=1}^{C}KL(f_{s}(\tilde{x}_{i};\tau^{s}),f_{t}^{{}^{\prime}}({x}_{i};\tilde{\tau}_{i}^{t}))$
$+(1-\alpha)\frac{1}{C}\sum_{i=1}^{C}KL(f_{s}({x}_{i};\tau^{s}),f_{t}^{{}^{\prime}}({x}_{i};\tau_{i}^{t})),$ (12)
where $KL$ represents Kullback–Leibler divergence loss, $\alpha$ is the trade-off parameter between accuracy and robustness. Meanwhile, we apply the anti-bias soft labels to generate adversarial examples $\tilde{x}$ as follows:
@@
In addition, to maintain stability when adjusting the teacher prediction distribution, we hold the student temperature $\tau^{s}$ constant, and the student’s optimization error risk for each class can be compared under the same standard.
@@ 5.1 Experimental Settings
We conduct our experiments on two benchmark datasets including CIFAR-10 [16] and CIFAR-100.
@@
Baselines. We consider the standard training method and eight state-of-the-art methods as comparison methods: Adversarial Training methods: SAT [18], and TRADES [39]; Adversarial Robustness Distillation methods: RSLAD [43], and AdaAD [12]; Class-wise Robust Fairness methods: FRL [37], BAT [26], CFA [34], and Fair-ARD [36].
@@
Student and Teacher Networks. For the student model, here we consider two networks for CIFAR-10 and CIFAR-100 including ResNet-18 [10] and MobileNet-v2 [24]. For the teacher model, we follow the setting in [43], and we select WiderResNet-34-10 [38] trained by [39] for CIFAR-10 and WiderResNet-70-16 trained by [9] for CIFAR-100.
Training Setting. For the whole baseline methods, we strictly follow their original setting if without additional instruction. For FRL [37], we use the Reweight+Remargin under the threshold of 0.05. For CFA [34], we select the best version (TRADES+CFA) as reported in the original article. For the training of CFA on CIFAR-100, we set the fairness threshold to 0.02 for FAWA operation based on the worst-10% class robustness. For Fair-ARD [36], we select the best version (Fair-ARD on CIFAR-10 and Fair-RSLAD on CIFAR-100) as reported in the original article.
Table 1: ResNet-18 robustness (%) on CIFAR-10 datasets. All results are the best checkpoint based on the best overall performance (Overall), which achieves the highest mean value of all-class average robustness (Avg.) and the worst class robustness (Worst).
Method Clean FGSM PGD ${}_{\scriptscriptstyle\rm{\infty}}$ AA
Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall
Natural 94.57 86.30 90.43 18.60 9.00 13.80 0 0 0 0 0 0 0 0 0
SAT[18] 84.03 63.90 73.96 56.71 26.70 41.70 49.34 21.00 35.17 48.99 19.90 34.44 46.44 16.80 31.62
TRADES[39] 81.45 67.60 74.52 56.65 36.60 46.62 51.78 30.40 41.09 49.15 27.10 38.12 48.17 25.90 37.03
RSLAD[43] 82.94 66.30 74.62 59.51 34.70 47.10 54.00 28.50 41.25 52.51 27.00 39.76 51.25 25.50 38.38
AdaAD[12] 84.73 68.10 76.41 59.70 34.80 47.25 53.82 29.30 41.56 52.30 26.00 39.15 50.91 24.70 37.80
FRL[37] 82.29 64.60 73.44 55.03 37.10 46.06 49.05 31.70 40.37 47.88 30.40 39.14 46.54 28.10 37.32
BAT[26] 86.72 72.30 79.51 60.97 33.80 47.38 49.60 22.70 36.15 47.49 19.50 33.49 48.18 20.70 34.44
CFA[34] 78.64 63.60 71.12 57.95 36.80 47.37 54.42 33.30 43.86 50.91 27.50 39.20 50.37 26.70 38.53
Fair-ARD[36] 83.81 69.40 76.76 58.41 38.50 48.46 50.91 29.90 40.41 49.96 28.30 39.13 47.97 25.10 36.54
ABSLD 83.04 68.10 75.57 59.83 40.50 50.17 54.50 36.50 45.50 51.77 32.80 42.29 50.25 31.00 40.63
Table 2: MobileNet-v2 robustness (%) on CIFAR-10 datasets. All results are the best checkpoint based on the best overall performance (Overall), which achieves the highest mean value of all-class average robustness (Avg.) and the worst class robustness (Worst).
Method Clean FGSM PGD ${}_{\scriptscriptstyle\rm{\infty}}$ AA
Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall
Natural 93.35 85.20 89.27 12.21 0.90 6.55 0 0 0 0 0 0 0 0 0
SAT[18] 82.30 65.80 74.05 56.19 34.60 45.39 48.52 24.90 36.71 47.22 23.80 35.51 44.38 18.50 31.44
TRADES[39] 79.37 59.00 69.18 54.94 31.60 43.27 50.03 27.20 38.61 47.02 23.20 35.11 46.25 22.10 34.17
RSLAD[43] 82.96 66.30 74.63 59.84 34.30 47.07 53.88 28.20 41.04 52.28 25.70 38.99 50.67 23.90 37.28
AdaAD[12] 83.72 66.90 75.31 57.63 33.70 45.66 51.89 26.90 39.39 50.34 24.70 37.52 48.81 22.50 35.65
FRL[37] 81.02 70.50 75.76 53.84 40.50 47.17 47.71 35.00 41.35 44.96 30.90 37.93 43.53 28.40 35.96
BAT[26] 83.01 70.30 76.65 53.01 32.10 42.55 44.08 25.00 34.54 41.85 21.60 31.72 42.65 23.40 33.02
CFA[34] 80.34 64.60 72.47 56.45 32.20 44.32 52.34 28.10 40.22 48.62 23.20 35.91 50.68 23.90 37.29
Fair-ARD[36] 82.44 69.40 75.92 56.29 38.60 47.45 50.91 29.90 40.41 48.18 30.80 39.49 46.62 27.60 37.11
ABSLD 82.54 69.50 76.02 58.55 41.40 49.98 52.99 35.70 44.35 50.39 31.90 41.15 48.71 30.30 39.51
Table 3: ResNet-18 robustness (%) on CIFAR-100 datasets. All results are the best checkpoint based on the best overall performance (Overall), which achieves the highest mean value of all-class average robustness (Avg.) and the worst-10% class robustness (Worst).
Method Clean FGSM PGD ${}_{\scriptscriptstyle\rm{\infty}}$ AA
Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall
Natural 75.17 53.80 64.48 7.95 0 3.97 0 0 0 0 0 0 0 0 0
SAT[18] 57.18 29.30 43.24 28.95 5.60 17.27 24.56 3.20 13.88 23.78 3.10 13.44 21.78 2.30 12.04
TRADES[39] 55.33 28.40 41.86 30.50 7.30 18.90 27.71 5.60 16.65 24.33 3.50 13.91 23.55 3.20 13.37
RSLAD[43] 57.88 29.40 43.64 34.50 9.20 21.85 31.19 7.40 19.29 28.13 4.70 16.41 26.46 3.70 15.08
AdaAD[12] 58.17 29.20 43.68 34.29 8.50 21.39 30.49 6.60 18.54 28.31 5.00 16.65 26.60 4.20 15.40
FRL[37] 55.49 30.30 42.89 28.00 7.20 17.60 24.04 5.00 14.52 22.93 3.70 13.31 21.10 2.80 11.95
BAT[26] 62.71 34.40 48.55 33.41 7.90 20.65 28.39 5.30 16.84 23.91 3.00 13.45 22.56 2.50 12.53
CFA[34] 59.29 33.80 46.54 33.88 10.30 22.09 31.15 8.40 19.77 26.85 5.20 16.02 25.83 4.90 15.36
Fair-ARD[36] 57.19 29.40 43.30 34.06 8.40 21.23 30.57 7.00 18.79 27.96 4.50 16.23 26.15 3.90 15.03
ABSLD 56.76 31.90 44.33 34.93 12.40 23.67 32.44 10.90 21.47 26.99 6.40 17.00 25.39 5.60 15.50
Table 4: MobileNet-v2 robustness (%) on CIFAR-100 datasets. All results are the best checkpoint based on the best overall performance (Overall), which achieves the highest mean value of all-class average robustness (Avg.) and the worst-10% class robustness (Worst).
Method Clean FGSM PGD ${}_{\scriptscriptstyle\rm{\infty}}$ AA
Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall Avg. Worst Overall
Natural 74.86 53.90 64.38 5.95 0 2.97 0 0 0 0 0 0 0 0 0
SAT[18] 56.70 22.90 39.80 32.10 7.00 19.55 28.61 5.30 16.95 26.55 3.50 15.02 24.36 2.40 13.38
TRADES[39] 57.10 30.00 43.55 31.70 8.20 19.95 29.43 6.20 17.81 25.25 3.70 14.47 24.39 3.20 13.79
RSLAD[43] 58.71 27.10 42.90 34.30 8.30 21.30 30.58 6.30 18.44 28.11 4.30 16.20 26.32 3.30 14.81
AdaAD[12] 54.59 24.50 39.54 31.40 6.10 18.75 27.96 4.80 16.38 25.72 2.10 13.91 23.80 1.60 12.70
FRL[37] 56.30 26.50 41.40 31.03 8.00 19.51 27.52 6.10 16.81 25.40 3.70 14.55 23.28 2.70 12.99
BAT[26] 65.39 38.70 52.04 33.31 5.30 19.30 27.08 2.90 14.99 22.80 1.50 12.15 21.28 1.10 11.19
CFA[34] 59.15 34.10 46.62 32.50 8.80 20.65 29.38 7.00 18.19 25.16 4.00 14.58 23.78 3.20 13.49
Fair-ARD[36] 58.97 31.40 45.19 33.76 9.30 21.53 30.07 7.30 18.69 27.64 5.60 16.62 25.79 4.00 14.90
ABSLD 56.66 32.00 44.33 33.87 12.80 23.34 31.24 11.40 21.32 26.41 7.50 16.96 24.57 6.70 15.64
For our ABSLD, we train the student using the Stochastic Gradient Descent (SGD) optimizer with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 2e-4. The learning rate $\beta$ of temperature is initially set as 0.1. For CIFAR-10 and CIFAR-100, we set the total number of training epochs to 300. The learning rate is divided by 10 at the 215-th, 260-th, and 285-th epochs; We set the batch size to 128 for both CIFAR-10 and CIFAR-100, which follows RSLAD [43]; meanwhile, the teacher temperature of $\tau_{k}^{t}$ and $\tilde{\tau}_{k}^{t}$ are initially set as 1 for all the classes. The student temperature $\tau_{s}$ is set to constant 1 without additional adjustment. With additional instruction, the maximum and minimum values of temperature are 5 and 0.5, respectively; For ResNet-18 on CIFAR-100, the maximum and minimum values of temperature are 3 and 0.8, respectively. For the inner maximization, we use a 10-step PGD with a random start size of 0.001 and a step size of 2/255. All training perturbation in the maximization is bounded to the $L_{\infty}$ norm $\epsilon$ = 8/255. For the training of the natural model, we train the models for 100 epochs, and the learning rate is divided by 10 at the 75-th and 90-th epochs.
Metrics. We evaluate the robust accuracy both in the average and the worst classes. For CIFAR-10, we report the worst class robust accuracy; For CIFAR-100, due to the poor performance of the worst class robustness, and only 100 images for each class in the test set, we follow the operation in CFA [34] and report the worst 10% class robust accuracy. Following previous studies [43, 41, 36], we evaluate the trained model against white box adversarial attacks: FGSM [8], PGD [18], ${}_{\scriptscriptstyle\rm{\infty}}$ [3]. For the PGD, we apply the 20 steps with a step size of 2/255; For ${}_{\scriptscriptstyle\rm{\infty}}$, we apply the 30 steps with a step size of 2/255. Meanwhile, we apply a strong attack: AutoAttack (AA) [5] to evaluate the robustness, which includes four attacks: Auto-PGD (APGD), Difference of Logits Ratio (DLR) attack, FAB-Attack [4], and the black-box Square Attack [1]. The maximum perturbation of all generated adversarial examples is 8/255.
Furthermore, as discussed in [34], the worst class robust accuracy changes drastically, so we follow [34] and select the best overall checkpoint, which achieves the highest mean value of all-class average robustness and the worst class robustness (where the worst class for CIFAR-10 and the worst 10% class for CIFAR-100) in all baselines and our method for a fair comparison. Meanwhile, we report the overall performance as a metric to evaluate the comprehensive performance for class-wise robust fairness.
5.2 Robustness and Fairness Performance
The performances of ResNet-18 and MobileNet-v2 trained by our ABSLD and other baseline methods under the various attacks are shown in Table 1, Table 2 for CIFAR-10, and Table 3, Table 4 for CIFAR-100. We select the best overall checkpoint for the baseline models and ABSLD.
Figure 3: The class-wise robustness of models guided by RSLAD and ABSLD on CIFAR-10, respectively. The robust accuracy is evaluated based on PGD. We can see that the harder classes’ robustness (class 3, 4, 5, 6) of ABSLD (blue lines) have different levels of improvement compared with RSLAD (red lines).
The results in Table 1, Table 2, Table 3, and Table 4 demonstrate that ABSLD achieves state-of-the-art robust fairness and overall performance on CIFAR-10 and CIFAR-100. For ResNet-18 on CIFAR-10, ABSLD improves the overall performance by 1.71%, 1.64%, 2.53%, and 2.10% compared with the best baseline method against the FGSM, PGD, ${}_{\scriptscriptstyle\rm{\infty}}$, and AA; For MobileNet-v2 on CIFAR-10, ABSLD improves the overall performance by 2.53%, 3.00%, 1.66%, and 2.22% compared with the best baseline method against the FGSM, PGD, ${}_{\scriptscriptstyle\rm{\infty}}$, and AA. Meanwhile, we visualize the class-wise robust performance of RSLAD and ABSLD in Figure 3, and the result shows that the robust performance of harder classes (class 3, 4, 5, 6) have different levels of improvement, which demonstrates that our method is beneficial to the overall robust fairness but not only to the worst class. Moreover, ABSLD shows relevant superiority on CIFAR-100 compared with other methods. The result shows that the robust fairness and overall performance of ABSLD has different levels of enhancement whether compared with the other adversarial training methods, adversarial robustness distillation methods, or class-wise robust fairness methods.

### v2

@@ 4.1 Overall Framework
Based on the above finding, adjusting the class-wise smoothness degree of soft labels can be regarded as a potential way to obtain robust fairness. Then we consider introducing this ideology into Knowledge Distillation (KD), which has been proven to be an effective method to improve the robustness of small models [44; 45; 43; 12; 42]. Since the core idea of KD is to use the teacher’s soft labels to guide the student’s optimization process, we can adjust the class-wise smoothness degree of soft labels and obtain the student with both strong robustness and fairness.
@@ 4.2 Re-temperate Teacher’s Soft Labels
In order to adjust the class-wise smoothness degree of soft labels in ARD, two options exist: one is to use student feedback to update the teacher parameter in the process of optimizing students, but this option requires retraining the teacher model, which may bring the pretty optimization difficulty and computational overhead; the other is to directly adjust the smoothness degree of soft labels for different classes. Inspired by [42], we apply the temperature as a means of directly controlling the smoothness degree of soft labels during the training process. Here, we provide Theorem 2 about the relationship between the teacher’s temperature and the student’s class-wise error risk gap.
Theorem 2.
@@
The proof of Theorem 2 can be found in Appendix A.2. In particular, just as the conclusion in [38]: The teacher has a more correct prediction distribution than the student even in the worst classes, which means Theorem 2 holds in most cases. Theorem 2 demonstrates that the different temperatures can influence the student robust fairness: when the student’s error risk of $k$-th class is larger than the average error risk, we think that this type of class is relatively hard compared with others, then the teacher temperature for $k$-th class will reduce and the smoothness degree of soft labels will be sharper, and the optimization gap between teacher distribution and student distribution in $k$-th class will corresponding increase, leading to stronger learning intensity for $k$-th class and final reduce the student’s class-wise optimization error risk gap.
@@
Meanwhile, according to [39], both clean and adversarial examples exist the fairness problems and can affect each other, so it is necessary to achieve fairness for both types of data. Since clean and adversarial examples of the same classes may have different error risks during the training process, it is unreasonable to use the same set of class temperatures to adjust both clean and adversarial examples. Here we simultaneously optimize the student’s clean error risk $R(f_{s}(x_{k}))$ and the student’s robust error risk $R(f_{s}(\tilde{x}_{k}))$, in other words, we apply two different sets of teacher temperatures: $\tau_{k}^{t}$ and $\tilde{\tau}_{k}^{t}$, for the adjustment of the teacher’s soft labels for clean and adversarial examples, respectively.
@@
1:the train dataset $\mathcal{D}$, Student $f_{s}$ with random initial weight $\theta_{s}$ and temperature $\tau_{s}$, pretrained robust teacher $f_{t}$, the initial temperature $\tau_{y}^{t}$ and $\tilde{\tau}_{y}^{t}$ for the teacher’s soft labels for clean examples $x$ and adversarial examples $\tilde{x}$, where $y=\{1,\dots,C\}$, the max training epochs $max$-$epoch$.
@@
4:          $R(f_{s}(\tilde{x}_{k}))=0,~{}R(f_{s}(x_{k}))=0$.  //  Initialize the clean and robust error risk for each class.
@@
7:          $\tilde{x}=\mathop{argmax}\limits_{||\tilde{x}-x||\leq\epsilon}KL(f_{s}(\tilde{x};\tau^{s}),f_{t}^{{}^{\prime}}(x;\tilde{\tau}_{y}^{t}))$.   //  Get adversarial examples with teacher’s soft labels.
8:          $\theta_{s}=\theta_{s}-\eta\cdot\nabla_{\theta}\mathcal{L}_{absld}(\tilde{x},x;f_{s},f_{t}^{{}^{\prime}})$.   //  Update student weight $\theta_{s}$ with teacher’s soft labels.
9:          $R(f_{s}(\tilde{x}_{y}))=R(f_{s}(\tilde{x}_{y}))+CE(f_{s}(\tilde{x}),{y})$.   //  Calculate robust error risk for each class.
10:          $R(f_{s}(x_{y}))=R(f_{s}(x_{y}))+CE(f_{s}({x}),{y})$.   //  Calculate clean error risk for each class.
@@
13:          Update $\tilde{\tau}_{k}^{t}$ and ${\tau}_{k}^{t}$ based on Eq.(10).   //  Re-temperate teacher’s soft labels for $\tilde{x}$ and $x$.
@@
$\mathcal{L}_{absld}(\tilde{x},x;f_{s},f_{t}^{{}^{\prime}})=\alpha\frac{1}{C}\sum_{i=1}^{C}KL(f_{s}(\tilde{x}_{i};\tau^{s}),f_{t}^{{}^{\prime}}({x}_{i};\tilde{\tau}_{i}^{t}))+(1-\alpha)\frac{1}{C}\sum_{i=1}^{C}KL(f_{s}({x}_{i};\tau^{s}),f_{t}^{{}^{\prime}}({x}_{i};\tau_{i}^{t})),$ (11)
where $KL$ represents Kullback–Leibler divergence loss, $\alpha$ is the trade-off parameter between accuracy and robustness, $f(x;\tau)$ denotes model $f$ predicts the output probability of $x$ with temperature $\tau$ in the final softmax layer. It should be mentioned that the teacher is frozen and we apply the teacher’s predicted soft labels $f_{t}^{{}^{\prime}}(x_{k};\tilde{\tau}_{k}^{t})$ for $k$-th class to generate adversarial examples $\tilde{x}_{k}$ as follows:
@@ 5.1 Experimental Settings
We conduct our experiments on three datasets: CIFAR-10 [16], CIFAR-100, and Tiny-ImageNet [17]. The results about CIFAR-100 and Tiny-ImageNet are in Appendix A.4 and A.5, respectively.
@@
Baselines. We consider the standard training method and eight state-of-the-art methods as comparison methods: AT methods: SAT [20], and TRADES [41]; ARD methods: RSLAD [45], and AdaAD [12]; Robust Fairness methods: FRL [39], BAT [28], CFA [36], and Fair-ARD [38].
@@
Student and Teacher Networks. For the student model, here we consider two networks for CIFAR-10 and CIFAR-100 including ResNet-18 [10] and MobileNet-v2 [26]. For the teacher model, we follow the setting in [45], and we select WiderResNet-34-10 [40] trained by [41] for CIFAR-10 and WiderResNet-70-16 trained by [9] for CIFAR-100. The teacher’s performance is in Appendix A.7.
@@
Method Clean FGSM PGD CW AA
Avg. Worst NSD Avg. Worst NSD Avg. Worst NSD Avg. Worst NSD Avg. Worst NSD
Natural 94.57 86.30 0.035 18.60 9.00 0.436 0 0 - 0 0 - 0 0 -
SAT[20] 84.03 63.90 0.118 56.71 26.70 0.283 49.34 21.00 0.332 48.99 19.90 0.352 46.44 16.80 0.385
TRADES[41] 81.45 67.60 0.113 56.65 36.60 0.267 51.78 30.40 0.301 49.15 27.10 0.341 48.17 25.90 0.350
RSLAD[45] 82.94 66.30 0.122 59.51 34.70 0.244 54.00 28.50 0.276 52.51 27.00 0.296 51.25 25.50 0.304
AdaAD[12] 84.73 68.10 0.114 59.70 34.80 0.246 53.82 29.30 0.285 52.30 26.00 0.312 50.91 24.70 0.322
[…]
BAT[28] 86.72 72.30 0.092 60.97 33.80 0.255 49.60 22.70 0.325 47.49 19.50 0.354 48.18 20.70 0.341
CFA[36] 78.64 63.60 0.123 57.95 36.80 0.231 54.42 33.30 0.258 50.91 27.50 0.288 50.37 26.70 0.296
Fair-ARD[38] 83.81 69.40 0.112 58.41 38.50 0.251 50.91 29.90 0.296 49.96 28.30 0.312 47.97 25.10 0.338
ABSLD 83.04 68.10 0.103 59.83 40.50 0.202 54.50 36.50 0.216 51.77 32.80 0.249 50.25 31.00 0.256
[…]
Method Clean FGSM PGD CW AA
Avg. Worst NSD Avg. Worst NSD Avg. Worst NSD Avg. Worst NSD Avg. Worst NSD
Natural 93.35 85.20 0.036 12.21 0.90 0.750 0 0 - 0 0 - 0 0 -
SAT[20] 82.30 65.80 0.131 56.19 34.60 0.279 48.52 24.90 0.334 47.22 23.80 0.361 44.38 18.50 0.410
TRADES[41] 79.37 59.00 0.131 54.94 31.60 0.297 50.03 27.20 0.330 47.02 23.20 0.367 46.25 22.10 0.379
RSLAD[45] 82.96 66.30 0.130 59.84 34.30 0.256 53.88 28.20 0.288 52.28 25.70 0.316 50.67 23.90 0.333
AdaAD[12] 83.72 66.90 0.123 57.63 33.70 0.262 51.89 26.90 0.300 50.34 24.70 0.317 48.81 22.50 0.334
[…]
BAT[28] 83.01 70.30 0.102 53.01 32.10 0.284 44.08 25.00 0.344 41.85 21.60 0.397 42.65 23.40 0.369
CFA[36] 80.34 64.60 0.112 56.45 32.20 0.260 52.34 28.10 0.292 48.62 23.20 0.320 50.68 23.90 0.331
Fair-ARD[38] 82.44 69.40 0.100 56.29 38.60 0.226 50.91 29.90 0.263 48.18 30.80 0.286 46.62 27.60 0.302
ABSLD 82.54 69.50 0.102 58.55 41.40 0.207 52.99 35.70 0.224 50.39 31.90 0.254 48.71 30.30 0.259
Training Setting. For ABSLD, we train the model using the Stochastic Gradient Descent (SGD) optimizer with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 2e-4. The learning rate $\beta$ of temperature is initially set as 0.1. For CIFAR-10 and CIFAR-100, we set the training epochs to 300. The learning rate is divided by 10 at the 215-th, 260-th, and 285-th epochs; We set the batch size to 128 for both CIFAR-10 and CIFAR-100 following [45]. For the inner maximization, we use a 10-step PGD with a random start size of 0.001 and a step size of 2/255, and the perturbation is bounded to the $L_{\infty}$ norm $\epsilon$ = 8/255. The more training setting can be found in Appendix A.3.
Metrics. We apply two metrics to evaluate the robust fairness: Normalized Standard Deviation (NSD [^2] NSD is a metric applied in [38] to measure the robust fairness. NSD = SD/Avg., where SD is the Standard Deviation of class-wise robustness and Avg. is the average robustness.) [38] and the worst-class robustness [36]. NSD can reflect robust fairness while also considering the average robustness. The smaller standard deviation means better fairness, and the larger average means better robustness, so the smaller NSD means better comprehensive performance in terms of fairness and robustness. The worst-class robustness is easy to understand, and a larger value means better fairness. For CIFAR-10, we directly report the worst class robust accuracy; For CIFAR-100 and Tiny-ImageNet, due to the poor performance of the worst class robustness and only 100 images (CIFAR-100) or 50 images (Tiny-ImageNet) for each class in the test set, we follow the operation in CFA [36] and report the worst 10% class robust accuracy. Besides, we also report the average robustness as a reference. The attack setting for evaluation can be found in Appendix A.3.
5.2 Robust Fairness Performance
The performances of ResNet-18 and MobileNet-v2 trained by our ABSLD and other baseline methods under the various attacks are shown in Table 1, Table 2 for CIFAR-10.
(a) Class-wise Robustness of ResNet-18 on CIFAR-10.
(b) Class-wise Robustness of MobileNet-v2 on CIFAR-10.
Figure 3: The class-wise robustness (PGD) of models guided by RSLAD and ABSLD on CIFAR-10. We can see that the harder classes’ robustness (class 3, 4, 5, 6) of ABSLD (blue lines) have different levels of improvement compared with RSLAD (red lines).
The results demonstrate that ABSLD achieves the state-of-the-art worst-class robustness on CIFAR-10. For ResNet-18 on CIFAR-10, ABSLD improves the worst class robustness by 2.0%, 3.2%, 2.4%, and 2.9% compared with the best baseline method against the FGSM, PGD, CW, and AA. Moreover, ABSLD shows relevant superiority on MobileNet-v2 compared with other methods.
Moreover, ABSLD can also show the best comprehensive performance of fairness and robustness (NSD) on CIFAR-10. For ResNet-18 on CIFAR-10, ABSLD reduces the NSD by 0.028, 0.032, 0.017, and 0.024 compared with the best baseline method against the FGSM, PGD, CW, and AA. The result indicates that although the trade-off between robustness and fairness still exists as [19] say, we obtain the highest robust fairness while sacrificing the least average robustness.
Meanwhile, we visualize the class-wise robustness in Figure 3, and the result shows that the robustness of harder classes (class 3, 4, 5, 6) have different levels of improvement, which demonstrates that our method is beneficial to the overall robust fairness but not only to the worst class. Moreover, combined with Figure 2, we can find that the trend of class-wise bias is similar in different training strategies, indicating that the bias may be sourced from the dataset itself, which further confirms Corollary 1.
In particular, we compare ABSLD with Fair-ARD [38], which is an adaptive re-weighting method on ARD. From the results, ABSLD has better robust fairness performance, which means that our proposed re-temperating method has superiority compared to the re-weighting method.
Figure 4: Ablation study for Baseline, Baseline+ABSLDadv, and ABSLD.
Figure 5: Standard deviation of class-wise clean optimization error risk.
Figure 6: Standard deviation of class-wise adversarial optimization error risk.